# Towards Explaining Deep Neural Network Compression Through a Probabilistic Latent Space

### v1

@@ 3 Theoretical Analysis
$D(F^{(L)},\widetilde{F}^{(L)}):=\Big{|}\mathbb{E}_{(\mathbf{X},Y)\sim D}\left\{L(F^{(L)}(\mathbf{X}),Y)\right\}-\mathbb{E}_{(\mathbf{X},Y)\sim D}\left\{L({\widetilde{F}}^{(L)}(\mathbf{X}),Y)\right\}\Big{|}.$ (7)
We are now in the position to state our first main finding and all proofs in details are provided in Appendix. Prior to introducing the theorems, it’s crucial to underline that each latent space is presupposed to include a covariance matrix among its inherent parameters.
3.1 Performance Difference Bounds
Theorem 1.
@@ 3.1 Performance Difference Bounds
where $g$ is an increasing function of $\epsilon^{(l)}$ i.e. for smaller $\epsilon^{(l)}$, the bound in (8) becomes tighter. Note that, trained sparse network reffers to the pruned network that has undergone fine-tuning after the pruning process.
@@
Theorem 1 is the immediate result of two Lemmas 2 and 3 below together. It is important to note that the function $g(\epsilon^{(l)})$ depends on $\epsilon^{(l)}$ sequentially through $\epsilon^{(L)},\ldots,\epsilon^{(l-1)}$. This means that as in Lemma 3, $\epsilon^{(l+1)}$ is an increasing function of $\epsilon^{(l)}$ for $l,\ldots,L-1$ and according to Lemma 2 the function $g$ is an increasing function of $\epsilon^{(L)}$.
In Lemma 2 and Corollary 2, $\lambda_{\text{max}}$ refers to the maximum eigenvalue of Hessian $\nabla^{2}\ell(\widetilde{\omega}^{*})$. Meanwhile, in Corollary 3, $\lambda_{\text{max}}$ represents the maximum eigenvalue of the inverse covariance matrix $\Sigma^{-1}$. In both corollaries, $\lambda_{\text{min}}$ denotes the minimum eigenvalue of $\Sigma^{-1}$. We also introduce constants $\epsilon^{(L)}$ and $\epsilon^{(l)}$ for reference.
Lemma 2.
Proof sketch of Lemma 2: While a detailed proof of Lemma 2 is available in Appendix, here’s a concise overview. By considering Equation 1 and by applying Jensen inequality, we have $D(F^{(L)},\widetilde{F}^{(L)})\leq\mathbb{E}_{(\mathbf{X},Y)\sim D}\left\{Y\cdot\left|\ell(\omega^{*})-\ell(\widetilde{\omega}^{*})\right|\right\}$. Using both quadratic Taylor approximation for $\ell$ and Corollary 1, $D(F^{(L)},\widetilde{F}^{(L)})\leq\mathbb{E}_{(\mathbf{X},Y)\sim D}\Big{\{}Y\cdot\Big{(}\frac{\lambda_{max}\sqrt{\frac{\epsilon^{(L)}}{2}}C}{2(1-\sqrt{\frac{\epsilon^{(L)}}{2}})}\Big{)}\Big{\}}$, where $C$ is a constant.
Corollary 2.
In Lemma 2, suppose the projection map function is Gaussian projection i.e. $P_{\omega^{*(L)}}\sim G(\omega^{*(L)},\Sigma)$ and $P_{\widetilde{\omega}^{*(L)}}\sim G(\widetilde{\omega}^{*(L)},\Sigma)$ with same positive definite covariance matrices. If the last layer of $F^{(L)}$ and $\widetilde{F}^{(L)}$ have AP3 with bound $\epsilon^{(L)}\geq 0$, then
$D(F^{(L)},\widetilde{F}^{(L)})\leq g_{G}^{(L)}(\epsilon^{(L)}),\;\;\hbox{where$g_{G}^{(L)}(\epsilon^{(L)})=\mathbb{E}_{(\mathbf{X},Y)\sim D}\Big{\{}Y\cdot\Big{(}\frac{\lambda_{max}}{2}\sqrt{\frac{2\epsilon^{(L)}}{\lambda_{min}}}\Big{)}\Big{\}}.$}$ (9)
Lemma 3.
Corollary 3.
In Lemma 3, suppose the projection map function is Gaussian projection i.e. $P_{\omega^{*(l)}}\sim G(\omega^{*(l)},\Sigma)$ and $P_{\widetilde{\omega}^{*(l)}}\sim G(\widetilde{\omega}^{*(l)},\Sigma)$ with covariance matrix $\Sigma$ for $l=~{}1,\ldots,L-1$. If for $\epsilon^{(l)}\geq 0$, $KL({P}_{\omega^{*(l)}}\|{P}_{\widetilde{\omega}^{*(l)}})\leq\epsilon^{(l)}$ and ${\rm Vol}(\mathcal{Z})<\infty$, then for constants $C_{1}$, $C_{\sigma^{-1},\sigma,x}$, and $K$,
$KL({P}_{\omega^{*(l+1)}}\|{P}_{\widetilde{\omega}^{*(l+1)}})\leq\epsilon^{(l+1)},\;\;\hbox{where \;$\epsilon^{(l+1)}:=\frac{1}{2}\left(C_{1}+\lambda_{max}C^{2}_{\sigma^{-1},\sigma,x}\big{(}\frac{2\epsilon^{(l)}-K}{\lambda_{min}}\big{)}\right)$}.$ (10)
3.2 How AP3 Explains Training Sparse Network?
@@
Theorem 1 links the AP3 property to the performance difference between Networks $F^{(L)}$ and $\widetilde{F}^{(L)}$ on optimal weights $\omega^{*}$ and $\widetilde{\omega}^{*}$. A follow-up question is how the latent space $\mathcal{Z}$ explains the training convergence of compressed network $\widetilde{F}^{(L)}$ to optimal sparsity level with approximately maintaining the performance. To provide an answer to this question we employ Definition 5 and approximately AP3 in iteration $t$ property as stated in the next theorem.
Theorem 2.
@@
Suppose that $P_{\omega^{*}}$ is the distribution of latent variable $Z\in\mathcal{Z}$ for network $F^{(L)}$ and $P_{\widetilde{\omega}^{*}}$ is the distribution of $\widetilde{Z}\in\widetilde{\mathcal{Z}}$ for sparse network $\widetilde{F}^{(L)}$. Let $\widetilde{\omega}^{t}$ be the weight of pruned network $\widetilde{F}_{t}^{(L)}$ after iteration $t$ of training with total $T$ iterations. Under the assumption that $F^{(L)}$ and $\widetilde{F}_{t}^{(L)}$ have approximately AP3 in iteration $t$ i.e. for small $\epsilon_{t}$, $|KL(P_{\omega^{*}}\|P_{\widetilde{\omega}^{*}})-KL(P_{\widetilde{\omega}^{*}}\|P_{\widetilde{\omega}^{t}})|=O(\epsilon_{t}),$ then we have $D({F}^{(L)},\widetilde{F}_{t}^{(L)})=O(\epsilon_{t},T)$, equivalently $\left|\mathbb{E}_{(\mathbf{X},Y)\sim D}\left[Y.\ell({\omega}^{*})\right]-\mathbb{E}_{(\mathbf{X},Y)\sim D}\left[Y.\ell(\widetilde{\omega}^{t})\right]\right|=O(\epsilon_{t},T)$.
4 Related Work
@@
Pruning has been a widely-used approach for reducing the size of a neural network by eliminating either unimportant weights or neurons. Existing methods can be categorized into structured and unstructured pruning. structured pruning, studied in [3, 6, 7, 8, 12, 13, 14, 15, 16], involves removing entire structures or components of a network, such as layers, channels, or filters. However, unstructured pruning aims to reduce the number of parameters by zeroing out some weights in the network. Some popular techniques commonly used in unstructured pruning are weight magnitude-based pruning, L1 regularization, Taylor expansion-based pruning, and Reinforcement learning-based pruning  [17, 4, 5, 18, 19, 20, 21].
@@
In this paper, to provide a comprehensive explanation of our findings, we focus on weight magnitude pruning technique, previously examined in [17, 19, 22]. In [17], where weight magnitude pruning is introduced, the authors suggested that a substantial number of network weights can be pruned without causing significant performance deterioration. To establish this notion, they conducted experiments employing models such as AlexNet and VGG16. While their research centers on empirically showcasing the efficacy of such technique, it doesn’t extensively explore the formal theoretical analysis of the subject. Our intent, instead, is to provide a theoretical foundation using probability latent spaces to elucidate the principles of network pruning, particularly weight pruning, and its significant role in preserving accuracy.
5 Experiments
@@
To validate our theoretical analysis empirically, contributing to a more comprehensive understanding of the sparsification techniques in DNNs, we conduct a set of experiments on two common CIFAR10 [23] and CIFAR100 [23] datasets, using three different pre-trained benchmark models: ResNet50 [24] and VGG16 [25], which are pre-trained on ImageNet. The results on AlexNet [26] are provided in Appendix.
5.1 Setup:
@@
Training Phase: We initialized these models’ weights with pre-trained weights as the baseline models. To ensure the results’ significance we performed three trials of training the models. Each trial involved training for 100 epochs using the SGD optimizer with a momentum of 0.9 and learning rates of 0.1 for pre-trained ResNet50 and 0.01 for AlexNet and VGG16. In order to facilitate convergence and enhance training efficiency, we incorporated a learning rate scheduler that dynamically adjusted the learning rate as the model approached convergence.
@@
Pruning Phase: We investigate the effects of weight differences within the last layer of the baseline and its pruned version, using magnitude-based pruning [17], over the PD. We initialize the model with the best parameters obtained from each trial. Three different pruning methods were employed: lowest (removing weights with the lowest magnitudes targeting connections that contributed the least to the model’s performance), highest (removing weights with the highest magnitudes, thus eliminating the most influential connections), and random (a stochastic element by randomly removing weights) magnitude prunings. Different percentages of sparsity level, including 0.1, 0.3, 0.5, and 0.8, are used to provide empirical evidence to support the validity of Lemma 2 for the optimal pruning percentage. We maintained similar experimental setups as before. However, given that we were fine-tuning the trained parameters after pruning, we limited the training to 20 epochs. Note that during each epoch of fine-tuning, the test accuracy of the pruned model is computed for the purpose of comparison with the test accuracy of the original network.
@@
The experimental results section is divided into three main segments, all centered around explaining optimal network sparsity using information-theoretic divergence measures: Section 5.2 investigates the relationship between AP2 and the performance difference of the original network and its sparse version by focusing on the last layer of the network, known as Lemma 2; Section 5.3 extends our analysis to AP3 by considering multivariate T-Student distribution; Finally, section 5.4 involves a crucial comparison between AP2 and AP3, aimed at determining whether AP2 truly constitutes the optimal explanation for PDs or if an opportunity exists to identify parameter sets within distributions that might surpass AP2’s capabilities.
5.2 Multivariate Gaussian Distribution with Diagonal Covariance Matrices
@@
In this section, we aim to validate Lemma 2 by examining the projections of the last layer weights of the networks and their pruned version onto multivariate Gaussian distributions. Specifically, we analyze the mean values of $\omega^{(L)}$ and $\widetilde{\omega}^{(L)}$ in their respective latent spaces, $\mathcal{Z}$ and $\mathcal{\widetilde{Z}}$. We utilize a coefficient of the identity covariance matrix for both latent spaces. By substituting this covariance matrix into the Gaussian distribution example outlined in Section 2, we observe that the KL-divergence value precisely matches AP2 (Gaussian Projection Example). Consequently, the following experiments provide strong evidence for the correctness of Lemma 2 for AP2 across various pre-trained models. Note that in Figs. 2 and 3 Y-axis shows the value of both AP2 and PD computed on testset.
@@
ResNet50: Fig. 2 shows a compelling visual demonstration, reaffirming the correctness of Lemma 2 through the use of AP2 on the pruned ResNet50 model applied to CIFAR10 and CIFAR100 datasets. It is crucial to emphasize that in Fig. 2, while the highest and random magnitude pruning methods consistently validate Lemma 2, the lowest magnitude pruning presents occasional challenges. Specifically, in cases where we prune $30\%$ and $50\%$ of the weights, the performance difference of these two pruning percentages is less than $10\%$. These discrepancies can be attributed to significant differences between theoretical assumptions and experimental observations. For instance, one crucial theoretical assumption in Lemma 2 is the use of optimal weights.
@@
Further, we conducted another experiment by comparing specific percentages of different pruning methods simultaneously, which is shown in the second row of Fig 2. Despite some variability due to stochastic gradient descent in certain epochs, the expected behavior described in Lemma 2 becomes evident. This signifies that when one pruning method’s KL-divergence decreases relative to another, its PD similarly diminishes. Note that covariance matrices for all methods and different percentages are the same in order to compare them in fairness.
@@
We ran additional experiments using Gaussian distribution with non-diagonal covariance matrices, which their KL-divergence serves as AP3. Detailed visual representations are included in Appendix to support our research findings.
5.3 Multivariate T-Student Distribution
Here, we employed the multivariate T-Student distribution, suited for heavy-tailed distributions and outliers in the latent space, enhancing the informativeness and interpretability of AP3 over AP2. We use the Monte Carlo technique to estimate KL-divergence in section 5.2. We set the degrees of freedom in the T-student distribution to $4$, aiming for heavier-tailed distributions and greater dissimilarity from a Gaussian distribution and generated 600,000 data points for improved estimation accuracy. To efficiently estimate KL divergence in high-dimensional spaces, we grouped weight vectors into 100 groups and averaged them, resulting in a 100-dimensional latent variable for KL-divergence estimation. While results for RexNet50 are presented here, experiments on AlexNet and VGG16 are detailed in Appendix. In Fig 4 (Y-axis shows both AP3 values and computed PD values on the test set), we observe that a reduction in KL-divergence of one method compared to another corresponds to a similar pattern in their PDs when meeting the assumptions in Lemma 2. However, some observations deviate from Lemma 2 due to AP3 values exceeding the threshold of $2$.
@@ 5.3 Multivariate T-Student Distribution
Figure 2: Comparing AP2 & PD for pruned ResNet50. Top- Comparing four pruning percentages for a given pruning method. Bottom-Comparing three pruning methods for a given pruning %. Expriments are done on CIFAR10 (Left) and CIFAR100 (Right).
This divergence is evident in the highest magnitude pruning of the first row of Fig. 4 (left) and $80\%$ lowest magnitude in the second row of Fig. 4 (right).
@@
Figure 3: Comparing AP2 & PD for pruned VGG16. Top- Comparing four pruning percentages for a given pruning method. Bottom-Comparing three pruning methods for a given pruning %. Expriments are done on CIFAR10 (Left) and CIFAR100 (Right).

### v3

@@ 2. Problem Formulation
(7) $D(F^{(L)},\widetilde{F}^{(L)}):=\Big{|}\mathbb{E}_{(\mathbf{X},Y)\sim D}\left\{L(F^{(L)}(\mathbf{X}),Y)\right\}-\mathbb{E}_{(\mathbf{X},Y)\sim D}\left\{L({\widetilde{F}}^{(L)}(\mathbf{X}),Y)\right\}\Big{|}.$
3. Theoretical Analysis
In DNNs compression, it is essential to maintain the performance with a core mission of reducing computational complexity and memory usage. However, in many compression methods, after significant computational improvement, we want the performance of the network after compression to be as close as possible to the best performance of the complete (non-compressed) network. In this section we state our main theoretical findings. All the proofs in detail are provided in the Appendix. Before introducing the theorems, it’s crucial to underline that each latent space is presupposed to include a covariance matrix among its inherent parameters.
3.1. Performance Difference Bounds
Lemma 3.1.
Proof sketch of Lemma 3.1: While a detailed proof of Lemma 3.1 is available in the Appendix, here’s a concise overview. By considering Equation 1 and by applying Jensen inequality, we have $D(F^{(L)},\widetilde{F}^{(L)})\leq\mathbb{E}_{(\mathbf{X},Y)\sim D}\left\{Y\cdot\left|\ell(\omega^{*})-\ell(\widetilde{\omega}^{*})\right|\right\}$. Using both quadratic Taylor approximation for $\ell$ and Corollary 2.6, $D(F^{(L)},\widetilde{F}^{(L)})\leq\mathbb{E}_{(\mathbf{X},Y)\sim D}\Big{\{}Y\cdot\Big{(}\frac{\lambda_{max}\sqrt{\frac{\epsilon^{(L)}}{2}}C}{2(1-\sqrt{\frac{\epsilon^{(L)}}{2}})}\Big{)}\Big{\}}$, where $C$ is a constant.
Corollary 3.2.
In Lemma 3.1, suppose the projection map function is Gaussian projection i.e. $P_{\omega^{*(L)}}\sim G(\omega^{*(L)},\Sigma)$ and $P_{\widetilde{\omega}^{*(L)}}\sim G(\widetilde{\omega}^{*(L)},\Sigma)$ with same positive definite covariance matrices. If the last layer of $F^{(L)}$ and $\widetilde{F}^{(L)}$ have AP3 with bound $\epsilon^{(L)}\geq 0$, then
(8) $D(F^{(L)},\widetilde{F}^{(L)})\leq g_{G}^{(L)}(\epsilon^{(L)}),\;\;\hbox{where$g_{G}^{(L)}(\epsilon^{(L)})=\mathbb{E}_{(\mathbf{X},Y)\sim D}\Big{\{}Y\cdot\Big{(}\frac{\lambda_{max}}{2}\sqrt{\frac{2\epsilon^{(L)}}{\lambda_{min}}}\Big{)}\Big{\}}.$}$
Lemma 3.3.
Corollary 3.4.
In Lemma 3.3, suppose the projection map function is Gaussian projection i.e. $P_{\omega^{*(l)}}\sim G(\omega^{*(l)},\Sigma)$ and $P_{\widetilde{\omega}^{*(l)}}\sim G(\widetilde{\omega}^{*(l)},\Sigma)$ with covariance matrix $\Sigma$ for $l=~{}1,\ldots,L-1$. If for $\epsilon^{(l)}\geq 0$, $KL({P}_{\omega^{*(l)}}\|{P}_{\widetilde{\omega}^{*(l)}})\leq\epsilon^{(l)}$ and ${\rm Vol}(\mathcal{Z})<\infty$, then for constants $C_{1}$, $C_{\sigma^{-1},\sigma,x}$, and $K$,
(9) $KL({P}_{\omega^{*(l+1)}}\|{P}_{\widetilde{\omega}^{*(l+1)}})\leq\epsilon^{(l+1)},\;\;\hbox{where \;$\epsilon^{(l+1)}:=\frac{1}{2}\left(C_{1}+\lambda_{max}C^{2}_{\sigma^{-1},\sigma,x}\big{(}\frac{2\epsilon^{(l)}-K}{\lambda_{min}}\big{)}\right)$}.$
In Lemma 3.1 and Corollary 3.2, $\lambda_{\text{max}}$ refers to the maximum eigenvalue of Hessian $\nabla^{2}\ell(\widetilde{\omega}^{*})$. Meanwhile, in Corollary 3.4, $\lambda_{\text{max}}$ represents the maximum eigenvalue of the inverse covariance matrix $\Sigma^{-1}$. In both corollaries, $\lambda_{\text{min}}$ denotes the minimum eigenvalue of $\Sigma^{-1}$. We also introduce constants $\epsilon^{(L)}$ and $\epsilon^{(l)}$ for reference.
Theorem 3.5.
@@ 3.1. Performance Difference Bounds
where $g$ is an increasing function of $\epsilon^{(l)}$ i.e. for smaller $\epsilon^{(l)}$, the bound in (10) becomes tighter. Note that, a trained sparse network refers to the pruned network that has undergone fine-tuning after the pruning process.
@@
Theorem 3.5 is the immediate result of two Lemmas 3.1 and 3.3 together. It is important to note that the function $g(\epsilon^{(l)})$ depends on $\epsilon^{(l)}$ sequentially through $\epsilon^{(L)},\ldots,\epsilon^{(l-1)}$. This means that as in Lemma 3.3, $\epsilon^{(l+1)}$ is an increasing function of $\epsilon^{(l)}$ for $l,\ldots,L-1$ and according to Lemma 3.1 the function $g$ is an increasing function of $\epsilon^{(L)}$.
It is important to note that while we do not claim AP2 and AP3 are always small, our empirical results in Section 5 demonstrate that whenever either metric falls below $\epsilon$, the resulting performance difference (PD) is correspondingly small. These metrics are especially meaningful in practice, as at least one widely used pruning method—Low-Magnitude Pruning—naturally tends to produce low AP2 values.
3.2. How AP3 Explains Training Sparse Network?
@@
Theorem 3.5 links the AP3 property to the performance difference between Networks $F^{(L)}$ and $\widetilde{F}^{(L)}$ on optimal weights $\omega^{*}$ and $\widetilde{\omega}^{*}$. A follow-up question is how the latent space $\mathcal{Z}$ explains the training convergence of compressed network $\widetilde{F}^{(L)}$ to optimal sparsity level with approximately maintaining the performance. To provide an answer to this question we employ Definition 2.7 and approximately AP3 in iteration $t$ property as stated in the next theorem.
Theorem 3.6.
@@
Suppose that $P_{\omega^{*}}$ is the distribution of latent variable $Z\in\mathcal{Z}$ for network $F^{(L)}$ and $P_{\widetilde{\omega}^{*}}$ is the distribution of $\widetilde{Z}\in\widetilde{\mathcal{Z}}$ for sparse network $\widetilde{F}^{(L)}$. Let $\widetilde{\omega}^{t}$ be the weight of pruned network $\widetilde{F}_{t}^{(L)}$ after iteration $t$ of training with total $T$ iterations. Under the assumption that $F^{(L)}$ and $\widetilde{F}_{t}^{(L)}$ have approximately AP3 in iteration $t$ i.e. for small $\epsilon_{t}$, $|KL(P_{\omega^{*}}\|P_{\widetilde{\omega}^{*}})-KL(P_{\widetilde{\omega}^{*}}\|P_{\widetilde{\omega}^{t}})|=O(\epsilon_{t}),$ then we have $D({F}^{(L)},\widetilde{F}_{t}^{(L)})=O(\epsilon_{t},T)$, equivalently
$\left|\mathbb{E}_{(\mathbf{X},Y)\sim D}\left[Y.\ell({\omega}^{*})\right]-\mathbb{E}_{(\mathbf{X},Y)\sim D}\left[Y.\ell(\widetilde{\omega}^{t})\right]\right|=O(\epsilon_{t},T).$
4. Related Work
@@
Pruning has been a widely-used approach for reducing the size of a neural network by eliminating either unimportant weights or neurons. Existing methods can be categorized into structured and unstructured pruning. structured pruning, studied in (Ding et al., 2019a, b; Liu et al., 2021; You et al., 2019; Anwar et al., 2017; Wen et al., 2016; Li et al., 2022; He et al., 2017; Luo et al., 2017), involves removing entire structures or components of a network, such as layers, channels, or filters. However, unstructured pruning aims to reduce the number of parameters by zeroing out some weights in the network. Some popular techniques commonly used in unstructured pruning are weight magnitude-based pruning, L1 regularization, Taylor expansion-based pruning, and Reinforcement learning-based pruning  (Han et al., 2015; Lee et al., 2020; Park et al., 2020; Sanh et al., 2020; Molchanov et al., 2019b; Zhu and Gupta, 2018; Ye et al., 2018).
@@
In this paper, to provide a comprehensive explanation of our findings, we focus on the weight magnitude pruning technique, previously examined in (Han et al., 2015; Molchanov et al., 2019b, a). In (Han et al., 2015), where weight magnitude pruning is introduced, the authors suggested that a substantial number of network weights can be pruned without causing significant performance deterioration. To establish this notion, they conducted experiments employing models such as AlexNet and VGG16. While their research centers on empirically showcasing the efficacy of such a technique, it doesn’t extensively explore the formal theoretical analysis of the subject. Our intent, instead, is to provide a theoretical foundation using probability latent spaces to elucidate the principles of network pruning, particularly weight pruning, and its significant role in preserving accuracy.
5. Experiments
@@
To validate our theoretical analysis empirically, contributing to a more comprehensive understanding of the sparsification techniques in DNNs, we conduct a set of experiments on three datasets, including CIFAR10 (Krizhevsky et al., 2009) and CIFAR100 (Krizhevsky et al., 2009) and Tiny-ImageNet, using four different pre-trained benchmark models: AlexNet (Krizhevsky et al., 2012), ResNet50 (He et al., 2016), VGG16 (Simonyan and Zisserman, 2015) and ViT (Nguyen et al., 2024), which are pre-trained on ImageNet. The results on AlexNet (Krizhevsky et al., 2012) are provided in the Appendix.
5.1. Setup:
@@
Training Phase: We initialized these models’ weights with pre-trained weights as the baseline models. To ensure the results’ significance we performed three trials of training the models. Each trial involved training for 100 epochs using the SGD optimizer with a momentum of 0.9 and learning rates of 0.1 for pre-trained ResNet50, 0.01 for AlexNet and VGG16 and 0.002 for Vit model. Considering the Tiny-ImageNet dataset, the learning rates are 0.01, 0.001, 0.002, respectively. The learning rate is set to $0.001$ fo all networks when the dataset is Tiny-ImageNet. In order to facilitate convergence and enhance training efficiency, we incorporated a learning rate scheduler that dynamically adjusted the learning rate as the model approached convergence.
@@
Pruning Phase: We investigate the effects of weight differences within the last layer of the baseline and its pruned version, using magnitude-based pruning (Han et al., 2015), over the PD. We initialize the model with the best parameters obtained from each trial. Three different pruning methods were employed: lowest (removing weights with the lowest magnitudes targeting connections that contributed the least to the model’s performance), highest (removing weights with the highest magnitudes, thus eliminating the most influential connections), and random (a stochastic element by randomly removing weights) magnitude prunings. Different percentages of sparsity level, including 0.1, 0.3, 0.5, and 0.8, are used to provide empirical evidence to support the validity of Lemma 3.1 for the optimal pruning percentage. We maintained similar experimental setups as before. However, given that we were fine-tuning the trained parameters after pruning, we limited the training to 20 epochs. Note that during each epoch of fine-tuning, the test accuracy of the pruned model is computed for the purpose of comparison with the test accuracy of the original network.
@@
The experimental results section is divided into three main segments, all centered around explaining optimal network sparsity using information-theoretic divergence measures: Section 5.2 investigates the relationship between AP2 and the performance difference of the original network and its sparse version by focusing on the last layer of the network, known as Lemma 3.1; Section 5.3 extends our analysis to AP3 by considering multivariate T-Student distribution; Finally, section 5.4 involves a crucial comparison between AP2 and AP3, aimed at determining whether AP2 truly constitutes the optimal explanation for PDs or if an opportunity exists to identify parameter sets within distributions that might surpass AP2’s capabilities.
5.2. Multivariate Gaussian Distribution with Diagonal Covariance Matrices
@@
In this section, we aim to validate Lemma 3.1 by examining the projections of the last layer weights of the networks and their pruned version onto multivariate Gaussian distributions. Specifically, we analyze the mean values of $\omega^{(L)}$ and $\widetilde{\omega}^{(L)}$ in their respective latent spaces, $\mathcal{Z}$ and $\mathcal{\widetilde{Z}}$. We utilize a coefficient of the identity covariance matrix for both latent spaces. By substituting this covariance matrix into the Gaussian distribution example outlined in Section 2, we observe that the KL-divergence value precisely matches AP2 (Gaussian Projection Example). Consequently, the following experiments provide strong evidence for the correctness of Lemma 3.1 for AP2 across various pre-trained models. Note that in Figs. 2, 3 and 4, Y-axis shows the value of both AP2 and PD computed on testset.
@@
ResNet50: Fig. 2 shows a compelling visual demonstration, reaffirming the correctness of Lemma 3.1 through the use of AP2 on the pruned ResNet50 model applied to CIFAR10 and CIFAR100 datasets. It is crucial to emphasize that in Fig. 2, while the highest and random magnitude pruning methods consistently validate Lemma 3.1, the lowest magnitude pruning presents occasional challenges. Specifically, in cases where we prune $30\%$ and $50\%$ of the weights, the performance difference of these two pruning percentages is less than $10\%$. These discrepancies can be attributed to significant differences between theoretical assumptions and experimental observations. For instance, one crucial theoretical assumption in Lemma 3.1 is the use of optimal weights.
@@
Further, we conducted another experiment by comparing specific percentages of different pruning methods simultaneously, which is shown in the second row of Fig 2. Despite some variability due to stochastic gradient descent in certain epochs, the expected behavior described in Lemma 3.1 becomes evident. This signifies that when one pruning method’s KL-divergence decreases relative to another, its PD similarly diminishes. Note that covariance matrices for all methods and different percentages are the same in order to compare them in fairness.
@@
To further validate our theoretical findings, we extended our analysis to the Tiny-ImageNet dataset. Fig. 4 illustrates a comparison of AP2 and PD on Tiny-ImageNet using both ResNet50 and VGG16 architectures. These results corroborate the generalizability of our theories. It is crucial to emphasize that, in Fig 4, random and lowest magnitude pruning methods exhibit trends that align more closely with the theoretical predictions described in the paper. In contrast, the highest magnitude pruning deviates significantly, especially at higher pruning percentages. One key reason for this deviation is that the theory assumes optimal pruning, which the highest magnitude pruning may fail to approximate in practice. Another critical factor is the increased complexity of the Tiny-ImageNet dataset. Networks trained on Tiny-ImageNet must learn and preserve class-specific, higher-level features, which are often encoded in the high-magnitude weights. Removing these weights disproportionately disrupts the latent representations necessary for handling Tiny-ImageNet’s finer granularity and a larger number of classes, leading to significant performance degradation.
We ran additional experiments using Gaussian distribution with non-diagonal covariance matrices, which their KL-divergence serves as AP3. Detailed visual representations are included in the Appendix to support our research findings.
@@
Figure 2. Comparing AP2 & PD for pruned ResNet50. Top- Comparing four pruning percentages for a given pruning method. Bottom-Comparing three pruning methods for a given pruning %. Expriments are done on CIFAR10 (Left) and CIFAR100 (Right).
@@
Figure 3. Comparing AP2 & PD for pruned VGG16. Top- Comparing four pruning percentages for a given pruning method. Bottom-Comparing three pruning methods for a given pruning %. Expriments are done on CIFAR10 (Left) and CIFAR100 (Right).
5.3. Multivariate T-Student Distribution
Here, we employed the multivariate T-Student distribution, suited for heavy-tailed distributions and outliers in the latent space, enhancing the informativeness and interpretability of AP3 over AP2. We use the Monte Carlo technique to estimate KL-divergence in section 5.2. We set the degrees of freedom in the T-student distribution to $4$, aiming for heavier-tailed distributions and greater dissimilarity from a Gaussian distribution and generated 600,000 data points for improved estimation accuracy. To efficiently estimate KL divergence in high-dimensional spaces, we grouped weight vectors into 100 groups and averaged them, resulting in a 100-dimensional latent variable for KL-divergence estimation. While results for RexNet50 and ViT are presented here, experiments on AlexNet and VGG16 are detailed in the Appendix. In Figs 5, 6 (Y-axis shows both AP3 values and computed PD values on the test set), we observe that a reduction in KL-divergence of one method compared to another corresponds to a similar pattern in their PDs when meeting the assumptions in Lemma 3.1. However, some observations deviate from Lemma 3.1 due to the use of a non-optimal pruning method, as evidenced by the large performance differences observed with this approach. This divergence is evident in the highest magnitude pruning of two figures  5 and  6.